# Curating art exhibitions using machine learning

### v1

@@ 3.1 Self-contained text vectorisation
Text vectorisation transforms human language in the form of text into a series of numbers, of fixed or variable length, that represent that text in a machine-readable format (Mikolov et al., 2013). That way, we can transform the title and the overview text (concatenated together) into a numerical vector – a series of ordered numbers (and we choose this to be of fixed size to keep the model complexity under control and to have a fixed number of input nodes on our neural network). This will serve as the $x$ input in our machine learning approach, a neural network, as depicted in the conceptual toy model in Figure 1. As implied by the toy model diagram, we therefore also need an output $y$ variable. One approach would be to have the actual artworks and metadata to be a running text output, like what we have on the raw data as depicted in the Listing 1 above. However, this would introduce two hard-to-solve challenges: first, the output is of variable size (length), since an exhibition can have any number of associated artworks, and also the artworks can have any number of metadata items associated with them; and second it would not be possible to constrain the neural network to select only artworks from the available $484,956$ works of art in the Met Museum. The first issue could potentially be solved by a so-called Long Short-Term Memory (LSTM) neural network (Hochreiter and Schmidhuber, 1997). These are used, for example, in one human language to another human language translation (or computer programming language to another), and have the advantage of allowing for variable input/output length. [^10] This ability to achieve variable output length is reached by using a neural network architecture with feedback loops, called a recurrent architecture (Elman, 1990). The neural network outputs a token (or word) at a time, and then that output token is concatenated with the input and fed back into the network until a special end-of-sequence (EOS) token is produced. This allows the model to have a flexible output length. Nonetheless, even with a LSTM network, and with only a small number of training examples like we have, it would be very difficult to get it to output a consistent format like the one we have for our raw data, with all those fields and fixed format metadata. Furthermore, we really want to constrain our machine learning method to the range of Met Museum artworks only, and not to a hypothetical range of all world artworks. Therefore, we have devised a simple approach that consists of two steps, first in flattening the string values of the fields (“Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”) for the training set in the format as depicted in Listing 2, and second, in calculating, from the count of those values (per exhibition), the probability of each value showing in each exhibition, as depicted in the excerpt in the Listing 3.
@@
Figure 2: Neural network approach using self-contained text vectorisation and the (fixed-length) probability of finding a value on the exhibitions metadata fields. From left to right, the neural network takes a variable length plain text, i.e., the title and description of the exhibition, then it creates a vectorisation, which outputs a numeric vector, an embedding of that text. That vector has 256 dimensions, or numbers. Then it passes that vector to a one-dimensional average pooling layer, which averages those 256-dimensional vectors to smaller 64-dimensional vectors. That is passed to a so-called hidden layer, a dense layer of 256 nodes that has an activation function (in this case a ReLU function) – the ReLU (Rectified Linear Unit) function is defined as $\text{ReLU}(x)=\max(0,x)$, and outputs simply the input if it is positive, and zero otherwise. Finally, a second hidden layer with 8615 nodes or neurons and a linear activation function (in this case identity function) maps the results to the probabilities. In the diagram, $y_{\text{dim}}$ is the number of distinct values of the exhibitions artworks metadata fields (e.g., [“European Sculpture and Decorative Arts”, “The American Wing”, “Diego de Pesquera”, “1585”, “Sculpture”, …]). Therefore, $y_{\text{dim}}=8615$. The final outputs are probabilities of each one of those metadata fields (or what we call generalised tags) to show up in the list of artworks associated with that input title and description. We tested several activation layer functions, which decide the type of output. We also played with the number of hidden layer nodes, and settled on 256 nodes. The network/training parameters were: number of epochs = 2048, with the text vectorisation using max_tokens = 32768, output_sequence_length = 256, output_mode = "int", and standardize = "lower_and_strip_punctuation". For all runs we used batch_size = 16.
We run and test the model using several combinations made of different parameters and different modes of operation. The neural network we use is depicted, in schematic and compact fashion, in Figure 2. The data processing starts left to right: we feed in a title and description, the “$x$” in Listing 3, and that text is cleaned and broken up; transformed into a vector; then after several layers, it outputs probabilities of each generalised tag to show up, and we shall use those later to map the most probable artworks from our database.
@@
From analysing the above Figure 3 we can conclude two key insights. First, that the models seem to converge as the MSE value decreases with epoch, as training progresses. Second, that it also seems to show that after 150 epochs or so, we start to have a serious case of over-fitting, as we can see the loss on the validation set stops decreasing, so the model is simply memorising the data after that.
@@
We also calculated the percentage of predicted artworks that match actual artworks, a very difficult task for our models to achieve since there are $400,000$+ artworks to choose from. The results are depicted below in Figure 5.
The results show that the model is best at predicting the “Department” and “Classification” generalised tags, e.g., Department = “Asian Art”, “Photographs”, “Drawings and Prints”, “European Sculpture and Decorative Arts”, “Modern and Contemporary Art”, etc., and Classification = “Prints”, “Photographs”, “Paintings”, “Drawings”, “Creche”, etc. The model, however, is quite bad at getting the “Artist Display Name”, which is not surprising but it still disappointing. So, the model seems to focus more on the style and historical periods or traditions of the art exhibition. It is also quite disappointing that the model cannot achieve more than a $50\%$ matching percentage in any of the generalised tags at all. Finally, the percentage of matched artworks, as depicted in Figure 5, is less than $1\%$. Even taken into account that we are trying to match a few artworks across a $400,000$+ artworks database, this is still very small. Taken together, the over-fitting, the low matching percentage of the generalised tags and the low matching of artworks seem to indicate the model is weak. We shall show that it is indeed very weak, next, when we test the model out-of-sample.
We then tested the model out-of-sample and input the following title and description for a hypothetical exhibition. For this case, we used all the 236 exhibitions as training set, since we are doing out-of-sample prediction, there is no need to restrict ourselves by splitting the full dataset into a training and validation set, it would be wasteful. Notice also that the choice of the description was done on purpose since the word “impressionist” only shows up twice in a single exhibition, the Met Museum’s “Neo-Impressionism” [^11] This exhibition website page can be seen in https://www.metmuseum.org/exhibitions/listings/2001/paul-signac-circle. exhibition of 2001, with 7 paintings, and the sentence “still life” shows up 50 times within the generalised tags, well below the count for other generalised tags as we can see from Table 2. So, we are trying to make it difficult for the machine learning model to predict what generalised tags, and therefore what artworks to choose for our test input exhibition.
@@
We then sort the database in descending order of $\text{hit}_{j}$, and extract the top 16 artworks that represent the best matches. The results, using the full training set of 236 exhibitions with a training of 2048 epochs were as follows (results were as bad at 32 and 64 and 150 epochs too, before it starts to do over-fitting):
@@
Figure 6: Marble composite of images related to the artworks from the Met Museum corresponding to the run with epochs=2048 for the model training with the entire dataset, using as input self-contained embeddings and output per exhibition metadata occurrence frequencies. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API (The Metropolitan Museum of Art, 2025a).
@@ 3.2 OpenAI-based embedding vectorisation
In this second approach, we use some external help in the form of the state-of-the-art OpenAI embedding model text-embedding-3-large that can perform very high-quality text vectorisation (OpenAI, 2024). So, instead of relying exclusively on the text input (a total of $26,388$ words) and its statistics to create numeric vectors that hopefully will show some syntactic and semantic similarity and statistical analysis that can be used, we shall rely on the much larger training corpus of the OpenAI models, which are known to produce numeric vectors that encode a high-level of syntactic and semantic representation (see e.g., Keraghel et al. (2024); Giglietto (2024); Petukhova et al. (2025)). We apply this to the input text, the exhibition’s title and description (overview text) concatenation, replace the encoding that was done fully internally before, while keeping the rest of the neural network unchanged, as described in the previous methodology as depicted in Figure 2. The resulting network can be seen in Figure 7 below.
@@
One of the advantages of using the OpenAI embedding models to create numeric vectors from the input text is that these models use the so-called Transformer architecture (Vaswani et al., 2017). These models take into account not only the word or token frequency, but also the direction or order of the text (so “…will many…” is encoded differently from “…many will…”) and contextual information such as recognising and recording that the “he” in “…the direction of Picasso’s work changed, and he moved away from Cubist approaches…” is related to “Picasso”. These models have semantic encoding, recording not just the probability distribution function of words but also the meaning, context and direction and far-away relationships. Furthermore, these models have been trained in a vast corpus of textual data, and therefore have better vector similarity, among other advantages, for semantically similar word sequences, particularly in English language.
@@
The results seem to show also an improvement with respect to the first model, the base model with self-contained embeddings (depicted in Figure 5). The results jump from around $0.35\%$ to around $0.8\%$. In fact, when we increase the number of epochs to 2048, we get a convergence and saturation of this intersection value at around $1.34\%$. Although this value seems small, one has to think carefully how much larger this value is with respect to the value one obtains if we select randomly $k$ artworks from the full $N=$484,956$$ works of art dataset. If we draw randomly $k=44$ ($44.36$ is the average number of artworks across all our 236 exhibitions) artworks from the full dataset, the average count of artworks expected overlap or intersection is $k\times\frac{k}{N}=\frac{k^{2}}{N}=\frac{44^{2}}{484956}=\approx 0.00399$. So, the percentage overlap (count of overlap divided by the number of artworks in each exhibition) can be approximated by $k\times\frac{k}{k\times N}=\frac{k}{N}=\frac{44}{484956}=\approx 9.072988\times 10^{-5}\approx 0.00907\%$. So, for our current model we can get a accuracy 150 times or so better than random choice (even for our basic self-contained model we get an accuracy 70 better than random choice). The question now is, do these better results reflect themselves on better out-of-sample predictions, where these models should really count?
@@
Figure 12: Marble composite of images from the Met Museum corresponding to the run with epochs=128 for the model training with the entire dataset, using as input OpenAI embeddings and output per exhibition metadata occurrence frequencies. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API (The Metropolitan Museum of Art, 2025a). Non-available images are marked by their object id JPEG name.
@@
These results, in particular the second one, with a longer training (even if it over-fits in the training set), are more in line with our expectations that our model is now working better as intended. We can also see that from the marble composite depicted in Figure 13. The first element that seems to improve was that we now have a much tighter artwork creation interval, all artworks are within the period 1777–1890, a span of 113 years for the run with 128 epochs and 1881–1890, a span of only 9 years, for the longer run with 2048 epochs. Impressionism is considered to be a period of art history around 1860–1886 (Herbert, 1988; Gombrich, 2023), although this is not exactly fixed, obviously. Nonetheless, our results are close, even if our system seems to have an inclination to choose a lot of Post-Impressionists and Neo-Impressionists rather than actual from the first Impressionists, as we asked it to do by using as input “…First Impressionists in Paris…”. The second element that seems to also improve, is that now we have all artworks “Department” = “European Paintings” as opposed to “Asian Art”, as it should be, and “Classification” as “Paintings”, not as “Prints” or “Textiles-Woven”. The model also seems to make an effort to find instances of still life, as can be seen by the 4 matches in Figure 12 and the 5 matches in Figure 13 – not only that is a good percentage of matches, but increases with increasing number of training epochs, as it should. [^12] We note that we can examine freely all the images of all the artworks, even if we cannot display them in this article, by going to the URL https://www.metmuseum.org/art/collection/search/xxxxxx, where xxxxxx is the object_id. This way we can decide, by visual inspection, which one are instances of still life or not.
Figure 13: Marble composite of images from the Met Museum corresponding to the run with epochs=2048 for the model training with the entire dataset, using as input OpenAI embeddings and output per exhibition metadata occurrence frequencies. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API (The Metropolitan Museum of Art, 2025a). Non-available images are marked by their object id JPEG name.
The results are also interesting in the sense that only $5$ out of the predicted $16$ artworks, $31.25\%$ of them, are present in our full dataset of 236 exhibitions, meaning that the model seems to be able to select artworks outside the training set, and select from the full dataset of $400,000$+$ artworks in the Met Museum. This was one of the objectives of this work, to be able to generalise outside the given input exhibition dataset. Therefore, we seem to be going in the right direction by using more advanced embeddings. In the next section, we go one step further in this direction.
@@ 3.3 Direct neural network embedding output to embedding metadata of artworks
Given that the OpenAI embeddings improved the accuracy of our model, we decided to use it not just for the input but for both the inputs and outputs, i.e., the text/descriptions of the exhibitions and the corresponding concatenated generalised tags for all the artworks in each exhibition. [^13] We note that there are several ways to encode the output target values, the generalised tags. One alternative would be to encode each exhibition against each of the generalised tags present in each of the subgroups, or one-to-one even, in effect creating and unfolding more training examples. We decided not to complicate this model too much and simply try to map the encoding or embedding of the title description against the encoding or embedding or a single long string corresponding to all the concatenated generalised tags for all the artworks in each exhibition in the order they show up on the exhibition webpage and the Met Museum database. We also create the embeddings for all the Met Museum’s $484,956$ works of art, for their corresponding generalised tags or metadata and store those on a vector database (Andoni and Indyk, 2008). The reasons we use a vector database are: first, the embeddings data structure becomes too large for our computational resources/capacity available; second, it becomes too slow to search for the best exact nearest neighbours in an array in memory/disk; and third, that these vector databases are optimal for storing embeddings and provide already built-in mechanisms for nearest neighbour search (Iscen et al., 2017), which will be the approach we take in order to find the artworks (or metadata embeddings) corresponding to the input texts (or title/description embeddings). We then train the model and when that trained model is passed an input text (either from the training or validation sets or from a new out-of-sample test set), it creates an output embedding vector, for which we can then find the nearest neighbours on the vector database of all Met Museum’s artworks embeddings. Those will be the selected artworks for that corresponding exhibition. We use the Facebook AI Similarity Search (FAISS) library as described in Douze et al. (2024), in particular we use the Inverted File with Flat (IndexIVFFlat)vectors model, since the exact match approach could not fit the memory available on our modest computational resources. Below in Figure 14 we can find a schematic of the neural network for this kind of modelling.
@@
The marble image for these works of art is depicted below in Figure 18.
Figure 18: Marble composite of images from the Met Museum corresponding to the run with epochs = 2048 for the model training with the entire dataset, using as input OpenAI embeddings of the title/description of each exhibition and output the OpenAI embeddings of concatenated generalised tags for all artworks in each exhibition. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API (The Metropolitan Museum of Art, 2025a). Non-available images are marked by their object id JPEG name.
The results in the above table and displayed in Figure 18 seem to show an improvement. In particular, we now find some first impressionists, such as Claude Monet and Camille Pissarro. We also have a high percentage of intersection of predicted artworks versus actual artworks on the validation set, as depicted in Figure 17, with an end value at 2048 epochs of around 400 times larger than the one expected by random choice. In terms of finding instances of still life themes, we find that we get only two still life instances, and furthermore these are from a cubist, Jean Metzinger, a painting from 1917, and from a symbolist, Fernand Khnopff, with a painting from 1884, as can be seen in Figure 18. Finally, the time interval of the predicted artworks, 1850–1923, 73 years, is much larger than the length of the impressionist period. So, this does not seem like a conclusive improvement over the previous simpler model, the one in Section 3.2. We now attempt using the full-scale power of OpenAI LLM models in the next section.

### v2

@@ 3.1 Self-contained text vectorisation
Text vectorisation transforms human language in the form of text into a series of numbers, of fixed or variable length, that represent that text in a machine-readable format [Mikolov et al., 2013]. That way, we can transform the title and the overview text (concatenated together) into a numerical vector – a series of ordered numbers (and we choose this to be of fixed size to keep the model complexity under control and to have a fixed number of input nodes on our neural network). This will serve as the $x$ input in our machine learning approach, a neural network, as depicted in the conceptual toy model in Figure 1. As implied by the toy model diagram, we therefore also need an output $y$ variable. One approach would be to have the actual artworks and metadata to be a running text output, like what we have on the raw data as depicted in the Listing 1 above. However, this would introduce two hard-to-solve challenges: first, the output is of variable size (length), since an exhibition can have any number of associated artworks, and also the artworks can have any number of metadata items associated with them; and second it would not be possible to constrain the neural network to select only artworks from the available $484,956$ works of art in the Met Museum. The first issue could potentially be solved by a so-called Long Short-Term Memory (LSTM) neural network [Hochreiter and Schmidhuber, 1997]. These are used, for example, in one human language to another human language translation (or computer programming language to another), and have the advantage of allowing for variable input/output length. [^10] This ability to achieve variable output length is reached by using a neural network architecture with feedback loops, called a recurrent architecture [Elman, 1990]. The neural network outputs a token (or word) at a time, and then that output token is concatenated with the input and fed back into the network until a special end-of-sequence (EOS) token is produced. This allows the model to have a flexible output length. Nonetheless, even with a LSTM network, and with only a small number of training examples like we have, it would be very difficult to get it to output a consistent format like the one we have for our raw data, with all those fields and fixed format metadata. Furthermore, we really want to constrain our machine learning method to the range of Met Museum artworks only, and not to a hypothetical range of all world artworks. Therefore, we have devised a simple approach that consists of two steps, first in flattening the string values of the fields (“Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”) for the training set in the format as depicted in Listing 2, and second, in calculating, from the count of those values (per exhibition), the probability of each value showing in each exhibition, as depicted in the excerpt in the Listing 3.
@@
Figure 2: Neural network approach using self-contained text vectorisation and the (fixed-length) probability of finding a value on the exhibitions metadata fields. From left to right, the neural network takes a variable length plain text, i.e., the title and description of the exhibition, then it creates a vectorisation, which outputs a numeric vector, an embedding of that text. That vector has 256 dimensions, or numbers. Then it passes that vector to a one-dimensional average pooling layer, which averages those 256-dimensional vectors to smaller 64-dimensional vectors. That is passed to a so-called hidden layer, a dense layer of 256 nodes that has an activation function (in this case a ReLU function) – the ReLU (Rectified Linear Unit) function is defined as $\text{ReLU}(x)=\max(0,x)$, and outputs simply the input if it is positive, and zero otherwise. Finally, a second hidden layer with 8615 nodes or neurons and a linear activation function (in this case identity function) maps the results to the probabilities. In the diagram, $y_{\text{dim}}$ is the number of distinct values of the exhibitions’ artworks metadata fields (e.g., [“European Sculpture and Decorative Arts”, “The American Wing”, “Diego de Pesquera”, “1585”, “Sculpture”, …]). Therefore, $y_{\text{dim}}=8615$. The final outputs are probabilities of each one of those metadata fields (or what we call generalised tags) to show up in the list of artworks associated with that input title and description. We tested several activation layer functions, which decide the type of output. We also played with the number of hidden layer nodes, and settled on 256 nodes. The network/training parameters were: number of epochs = 2048, with the text vectorisation using max_tokens = 32768, output_sequence_length = 256, output_mode = "int", and standardize = "lower_and_strip_punctuation". For all runs we used batch_size = 16.
We run and test the model using several combinations made of different parameters and different modes of operation. The neural network we use is depicted, in schematic and compact fashion, in Figure 2. The data processing starts left to right: we feed in a title and description, the “xx” in Listing 3, and that text is cleaned and broken up; transformed into a vector; then after several layers, it outputs probabilities of each generalised tag to show up, and we shall use those later to map the most probable artworks from our database.
@@
From analysing Figure 3 we can conclude two key insights. First, that the models seem to converge as the MSE value decreases with epoch, as training progresses. Second, that it also seems to show that after 150 epochs or so, we start to have a serious case of over-fitting, as we can see the loss on the validation set stops decreasing, so the model is simply memorising the data after that.
@@
The results show that the model is best at predicting the “Department” and “Classification” generalised tags, e.g., Department = “Asian Art”, “Photographs”, “Drawings and Prints”, “European Sculpture and Decorative Arts”, “Modern and Contemporary Art”, etc., and Classification = “Prints”, “Photographs”, “Paintings”, “Drawings”, “Creche”, etc. The model, however, is quite bad at getting the “Artist Display Name”, which is not surprising but it still disappointing. So, the model seems to focus more on the style and historical periods or traditions of the art exhibition. It is also quite disappointing that the model cannot achieve more than a $50\%$ matching percentage in any of the generalised tags at all.
@@
We then sort the database in descending order of $\text{hit}_{j}$, and extract the top 16 artworks that represent the best matches. Given this, we can calculate the percentage of matched artworks, as depicted in Figure 5, is less than $1\%$. Even taken into account that we are trying to match a few artworks across a 400,000400,000+ artworks database, this is still very small. Taken together, the over-fitting, the low matching percentage of the generalised tags and the low matching of artworks seem to indicate the model is weak. We shall show that it is indeed very weak, next, when we test the model out-of-sample.
We tested the model out-of-sample and input the following title and description for a hypothetical exhibition. For this case, we used all the 236 exhibitions as training set, since we are doing out-of-sample prediction, there is no need to restrict ourselves by splitting the full dataset into a training and validation set, it would be wasteful. Notice also that the choice of the description was done on purpose since the word “impressionist” only shows up twice in a single exhibition, the Met Museum’s “Neo-Impressionism” [^11] This exhibition website page can be seen in https://www.metmuseum.org/exhibitions/listings/2001/paul-signac-circle. exhibition of 2001, with 7 paintings, and the sentence “still life” shows up 50 times within the generalised tags, well below the count for other generalised tags as we can see from Table 2. So, we are trying to make it difficult for the machine learning model to predict what generalised tags, and therefore what artworks to choose for our test input exhibition.
The results, using the full training set of 236 exhibitions with a training of 2048 epochs were as follows (results were as bad at 32 and 64 and 150 epochs too, before it starts to do over-fitting):
@@
Figure 6: Marble composite of images related to the artworks from the Met Museum corresponding to the run with epochs=2048 for the model training with the entire dataset, using as input self-contained embeddings and output per exhibition metadata occurrence frequencies. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API [The Metropolitan Museum of Art, 2025a].
@@ 3.2 OpenAI-based embedding vectorisation
In this second approach, we use some external help in the form of the state-of-the-art OpenAI embedding model text-embedding-3-large that can perform very high-quality text vectorisation [OpenAI, 2024]. So, instead of relying exclusively on the text input (a total of $26,388$ words) and its statistics to create numeric vectors that hopefully will show some syntactic and semantic similarity and statistical analysis that can be used, we shall rely on the much larger training corpus of the OpenAI models, which are known to produce numeric vectors that encode a high-level of syntactic and semantic representation (see e.g., Keraghel et al. [2024], Giglietto [2024], Petukhova et al. [2025]). We apply this to the input text, the exhibition’s title and description (overview text) concatenation, replace the encoding that was done fully internally before, while keeping the rest of the neural network unchanged, as described in the previous methodology as depicted in Figure 2. The resulting network can be seen in Figure 7 below.
@@
One of the advantages of using the OpenAI embedding models to create numeric vectors from the input text is that these models use the so-called Transformer architecture [Vaswani et al., 2017]. These models take into account not only the word or token frequency, but also the direction or order of the text (so “…will many…” is encoded differently from “…many will…”) and contextual information such as recognising and recording that the “he” in “…the direction of Picasso’s work changed, and he moved away from Cubist approaches…” is related to “Picasso”. These models have semantic encoding, recording not just the probability distribution function of words but also the meaning, context and direction and far-away relationships. Furthermore, these models have been trained in a vast corpus of textual data, and therefore have better vector similarity, among other advantages, for semantically similar word sequences, particularly in English language.
@@
The results seem to show also an improvement with respect to the first model, the base model with self-contained embeddings (depicted in Figure 5). The results jump from around $0.35\%$ to around $0.8\%$. In fact, when we increase the number of epochs to 2048, we get a convergence and saturation of this intersection value at around $1.34\%$. Although this value seems small, one has to think carefully how much larger this value is with respect to the value one obtains if we select randomly $k$ artworks from the full N=484,956N=\$484,956\$ works of art dataset. If we draw randomly $k=44$ ($44.36$ is the average number of artworks across all our 236 exhibitions) artworks from the full dataset, the average count of artworks expected overlap or intersection is $k\times\frac{k}{N}=\frac{k^{2}}{N}=\frac{44^{2}}{484956}=\approx 0.00399$. So, the percentage overlap (count of overlap divided by the number of artworks in each exhibition) can be approximated by $k\times\frac{k}{k\times N}=\frac{k}{N}=\frac{44}{484956}=\approx 9.072988\times 10^{-5}\approx 0.00907\%$. So, for our current model we can get a accuracy 150 times or so better than random choice (even for our basic self-contained model we get an accuracy 70 better than random choice). The question now is, do these better results reflect themselves on better out-of-sample predictions, where these models should really count?
@@
Figure 12: Marble composite of images from the Met Museum corresponding to the run with epochs=128 for the model training with the entire dataset, using as input OpenAI embeddings and output per exhibition metadata occurrence frequencies. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API [The Metropolitan Museum of Art, 2025a]. Non-available images are marked by their object id JPEG name.
@@
These results, in particular the second one, with a longer training (even if it over-fits in the training set), are more in line with our expectations that our model is now working better as intended. We can also see that from the marble composite depicted in Figure 13. The first element that seems to improve was that we now have a much tighter artwork creation interval, all artworks are within the period 1777–1890, a span of 113 years for the run with 128 epochs and 1881–1890, a span of only 9 years, for the longer run with 2048 epochs. Impressionism is considered to be a period of art history around 1860–1886 [Herbert, 1988, Gombrich, 2023], although this is not exactly fixed, obviously. Nonetheless, our results are close, even if our system seems to have an inclination to choose a lot of Post-Impressionists and Neo-Impressionists rather than actual from the first Impressionists, as we asked it to do by using as input “…First Impressionists in Paris…”. The second element that seems to also improve, is that now we have all artworks “Department” = “European Paintings” as opposed to “Asian Art”, as it should be, and “Classification” as “Paintings”, not as “Prints” or “Textiles-Woven”. The model also seems to make an effort to find instances of still life, as can be seen by the 4 matches in Figure 12 and the 5 matches in Figure 13 – not only that is a good percentage of matches, but increases with increasing number of training epochs, as it should. [^12] We note that we can examine freely all the images of all the artworks, even if we cannot display them in this article, by going to the URL https://www.metmuseum.org/art/collection/search/xxxxxx, where xxxxxx is the object_id. This way we can decide, by visual inspection, which one are instances of still life or not.
Figure 13: Marble composite of images from the Met Museum corresponding to the run with epochs=2048 for the model training with the entire dataset, using as input OpenAI embeddings and output per exhibition metadata occurrence frequencies. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API [The Metropolitan Museum of Art, 2025a]. Non-available images are marked by their object id JPEG name.
The results are also interesting in the sense that only $5$ out of the predicted $16$ artworks, $31.25\%$ of them, are present in our full dataset of 236 exhibitions, meaning that the model seems to be able to select artworks outside the training set, and select from the full dataset of 400,000+\$400,000\$+ artworks in the Met Museum. This was one of the objectives of this work, to be able to generalise outside the given input exhibition dataset. Therefore, we seem to be going in the right direction by using more advanced embeddings. In the next section, we go one step further in this direction.
@@ 3.3 Direct neural network embedding output to embedding metadata of artworks
Given that the OpenAI embeddings improved the accuracy of our model, we decided to use it not just for the input but for both the inputs and outputs, i.e., the text/descriptions of the exhibitions and the corresponding concatenated generalised tags for all the artworks in each exhibition. [^13] We note that there are several ways to encode the output target values, the generalised tags. One alternative would be to encode each exhibition against each of the generalised tags present in each of the subgroups, or one-to-one even, in effect creating and unfolding more training examples. We decided not to complicate this model too much and simply try to map the encoding or embedding of the title description against the encoding or embedding or a single long string corresponding to all the concatenated generalised tags for all the artworks in each exhibition in the order they show up on the exhibition webpage and the Met Museum database. We also create the embeddings for all the Met Museum’s $484,956$ works of art, for their corresponding generalised tags or metadata and store those on a vector database [Andoni and Indyk, 2008]. The reasons we use a vector database are: first, the embeddings data structure becomes too large for our computational resources/capacity available; second, it becomes too slow to search for the best exact nearest neighbours in an array in memory/disk; and third, that these vector databases are optimal for storing embeddings and provide already built-in mechanisms for nearest neighbour search [Iscen et al., 2017], which will be the approach we take in order to find the artworks (or metadata embeddings) corresponding to the input texts (or title/description embeddings). We then train the model and when that trained model is passed an input text (either from the training or validation sets or from a new out-of-sample test set), it creates an output embedding vector, for which we can then find the nearest neighbours on the vector database of all Met Museum’s artworks embeddings. Those will be the selected artworks for that corresponding exhibition. We use the Facebook AI Similarity Search (FAISS) library as described in Douze et al. [2024], in particular we use the Inverted File with Flat (IndexIVFFlat) vectors model, since the exact match approach could not fit the memory available on our modest computational resources. Below in Figure 14 we can find a schematic of the neural network for this kind of modelling.
@@
The marble image for these works of art is depicted below in Figure 18. The results in the above table and displayed in Figure 18 seem to show an improvement. In particular, we now find some early impressionists, such as Claude Monet and Camille Pissarro. We also have a high percentage of intersection of predicted artworks versus actual artworks on the validation set, as depicted in Figure 17, with an end value at 2048 epochs of around 400 times larger than the one expected by random choice. In terms of finding instances of still life themes, we find that we get only two still life instances, and furthermore these are from a cubist, Jean Metzinger, a painting from 1917, and from a symbolist, Fernand Khnopff, with a painting from 1884, as can be seen in Figure 18. Finally, the time interval of the predicted artworks, 1850–1923, 73 years, is much larger than the length of the impressionist period. So, this does not seem like a conclusive improvement over the previous simpler model, the one in Section 3.2. We now attempt using the full-scale power of OpenAI LLM models in the next section.
Figure 18: Marble composite of images from the Met Museum corresponding to the run with epochs = 2048 for the model training with the entire dataset, using as input OpenAI embeddings of the title/description of each exhibition and output the OpenAI embeddings of concatenated generalised tags for all artworks in each exhibition. Notice we only display the images of the artworks that are on the public domain, as provided by the Met Museum website and its API [The Metropolitan Museum of Art, 2025a]. Non-available images are marked by their object id JPEG name.